KAGRA information


CAL group corrected the calibration, the renewed binary range is now monitored in the control room.

(2)IFO status

(a) The longest lock was extended up to 6h 56m at present, and unlocked at 09:34.

(b)One SAL alignment with many troubles was applied yesterday noon time.

We found several problems in SAL alignment in Y arm an OMC dither. We should learn from Ushiba-kun what should we do during Y arm alignment. So IFO situation could not be same with on 7th April. Actually, the sensitivity has excess white like noise around 100-300Hz. We might have investigation break for this problem.

Please remind that SAL alignment is the final final method to recover the IFO, because LSC_LOCK guardian and observation mode control has a function to chase and record good alignment.

(c) Yuzurihara-kun prepared a task "noise investigation tool" for co-operator. The manual is here,


please co-operator try from now.

(d) Shift Summary are here,





mid night http://klog.icrr.u-tokyo.ac.jp/osl/?r=14053

General information from KAGRA office


KAGRA/Observation/KAGRA-1/ObservationApril7toApril21/9th (last edited 2020-04-10 13:44:03 by miyoki)